# Machine learning analysis of structural data to predict electronic properties in near-surface InAs quantum wells

### v1

@@ Abstract
Figure 1: a) Schematic of near-surface InAs quantum well heterostructure. b) Exemplary 20x20 $\mu$m AFM image for a near-surface InAs quantum well sample JS629. c) Schematic of crosshatch waveform model, adapted from [4]. d)-f) Plots of quantum well mobility as a function of the average crosshatch amplitude ($\overline{\epsilon}$) and wavelength ($\overline{\lambda}$).
um well hetero tructures within the last decade have been extensively studied for topological superconductivity [5, 6] and gate-tunable superconducting circuits for quantum information[7, 8, 9, 10, 11]. These studies utilize material stacks in which the quantum well region is placed close to the surface of the wafer. In this case, the active region of the wafer is well-separated from the initial mismatched interface by the growth of substantial metamorphic buffer layers [12, 13]. Compared to buried quantum well structures which are well known to have carrier scattering dominated by background and ionized impurities [14], modulation in the surface structure and surface dangling bonds have a much stronger impact on near-surface quantum well systems. As a result, understanding the surface strain state and strain fluctuations becomes of greater importance. Therefore, in these heterostructures it is of great significance to develop models with which we can quickly understand and predict how these materials will perform in cryogenic applications without the need for long fabrication, cooldown, and measurement cycles.
@@
Figure 2: Deconvolution of AFM image. a) Averaged color map for vertical crosshatch channel of sample JS629. b) and c) are the vertical and horizontal channels of sample JS629 after threshold processing to define the peaks and valleys of the waveform. d) Overlaid vertical and horizontal channels from our AFM image processes algorithm on top of the original rotated AFM image from Fig. 1b.
InAs near-surface quantum wells are grown in a Varian Gen II molecular beam epitaxy (MBE) system on 50.8 mm InP (001) wafers. The InAs quantum well structures are grown on a strain-relaxed graded buffer layer with top-layer composition of In0.81Al0.19As. The quantum well is grown to be 4 nm thick and is separated from the wafer surface by a 10 nm thick In0.81Ga0.19As barrier layer. More details and specifics about this structure growth can be found in reference [12]. After growth, AFM imaging is done with a NanoMagnetics Instruments AFM operating in a dynamical tapping mode. Preliminary AFM image normalization and processing is done using the Gwyddion software package [15]. Afterwards, quantum Hall transport is measured in an Oxford TelsatronPT 1.5 K He4 cryostat. The 2-in wafers are diced into roughly 5$\times$5 mm pieces which then are etched in Transene Type-D (Transene Company) aluminum etchant to remove the top Al layer. Samples are then cleaned in solvents, rinsed in de-ionized water, and then blow-dried with dry nitrogen. Measurements are done in a Van der Pauw wiring configuration.
Figure 3: Extraction of $\lambda$. Histogram plots in a) and b) for the vertical and horizontal channels, respectively, for the exemplary AFM image of sample JS629, presented in Fig. 1a. The overlaid line in both plots is a Gaussian fit to the histogram from which we extract a center of mass value and a line width. These values are listed in the legend of this plot.
@@
Figure 4: Linear regression model fitting. a) Linear regression model fitting to the full dataset. The red line is the model fit. Calculated R2 and p-values are stated in the inset. b) Linear regression model fitting with $\lambda_{x}$ and $\lambda_{y}$ removed from dataset. The red line is the model fit. c) Calculated residuals for model 1. d) Calculated residuals for model 2.
Once the AFM images are processed and we obtain the thresholded channel plots as shown in Figs. 2b and c, we then extract the widths of each line and gap between lines. These values are plotted in separate histogram plots for the vertical and horizontal channels, examples of which are shown in Figs. 3a and b. The histogram shows extracted values for the AFM image presented in Fig. 1b and corresponds to the de-convoluted crosshatch graphs presented in Fig. 2b and c. en fit this histogram to a Gaussian distribution to extract a center of mass (CoM) value and the full-width at half max (FWHM). In the context of a real material evaluation, the CoM is related to the crosshatch density (residual strain state), whereas the FWHM value is then related to the fluctuations in this strain state.
@@
We then fit the variables from the compiled dataset across all 36 samples to a multiple linear regression model using the stats package within base R statistical software (ver 4.3.3) [16]. The results of two of these model fits are presented in Fig. 4. Figs. 4a,c are the regression model and calculated residuals for Model 1 and Figs. 4b,d are the same for Model 2. The difference between models is that for Model 2, the $\overline{\lambda}_{x}$ and $\overline{\lambda}_{y}$ variables are removed from the fit. The resulting linear regression for Model 1 and 2 are presented in Equations 1 and 2, respectively.
$\begin{split}\overline{\lambda}_{MFP}&=60.59-0.01\overline{\lambda}_{x}-0.05\overline{\lambda}_{y}-35.98\overline{\epsilon}_{x}-50.01\overline{\epsilon}_{y}\\ &+0.39\sigma^{\lambda}_{x}-0.29\sigma^{\lambda}_{y}+169.72\sigma^{\epsilon}_{x}+35.59\sigma^{\epsilon}_{y}\end{split}$ (1)
$\begin{split}\overline{\lambda}_{MFP}&=57.76-33.45\overline{\epsilon}_{x}-49.2\overline{\epsilon}_{y}+0.4\sigma^{\lambda}_{x}-0.39\sigma^{\lambda}_{y}\\ &+168.02\sigma^{\epsilon}_{x}+15.59\sigma^{\epsilon}_{y}\end{split}$ (2)
The plots presented in Fig. 4a,b show the predicted (Model) values for electron mean free path, $\overline{\lambda}_{MFP}$, as a function of the experimental $\overline{\lambda}_{MFP}$ value. The model $\overline{\lambda}_{MFP}$ value is calculated using the linear regression line shown in Equations 1 and 2 for Figs. 4a and b, respectively, using the original experimental dataset that is obtained through our computer vision algorithm. The $R^{2}$ and p-values for each model fit are presented in the insets in Figs. 4a and b.
@@
Figure 5: Crosshatch surface modeling. a) Simulated crosshatch surface for a starting density of $1.4\times 10^{9}/cm^{2}$. b) and c) present exemplary experimental (“original”) and the 2D simulated AFM images, respectively, for comparison. The simulated image presented here is the 2D representation of the 3D crosshatch surface presented in a). d) and e): $\overline{\lambda}_{x}$ and $\overline{\lambda}_{y}$, respectively, as a function of dislocation density from our crosshatch simulations. f) and g): $\overline{\epsilon}_{x}$ and $\overline{\epsilon}_{y}$ as a function of dislocation density from our crosshatch simulations. In d)-g), the red lines represent the average value across the repeat simulations for each dislocation density while the grey shaded area represents the first standard deviation.
@@
Our AFM simulation utilizes the model for surface crosshatch proposed by Andrews et al. in 2002 [3]. The original model solves Hooke’s law in one dimension for an arbitrary elastic medium. The model is initialized by defining some starting density of dislocations along the a 1D line, and then integrating the strain field caused by the dislocations to a specified distance away from this starting point. This results in us obtaining an expression for the displacement field of the elastic medium at an arbitrary distance away. We first expanded the model to include two dimensions to generate a surface crosshatch in InAs from a given dislocation density. This model randomly places dislocations within a 20x20 $\mu$m region based on an initial user-defined dislocation density. The model equations are presented below in Equations 3 and 4 for calculating the displacement field in x and y, respectively.
@@
To generate the simulated dataset, for each dislocation density we simulate the surface 1.5 $\mu$m away from the misfit interface, consistent with the thickness of the material stack used in our experiments. For simplicity, we assume the majority of dislocations form at the interface with the InP substrate. The simulation was run for 1000 iterations per dislocation density, for 18 different densities ranging from $3\times 10^{7}/cm^{2}$ to $3\times 10^{9}/cm^{2}$. With this, we then extract the simulated $\lambda$ and $\epsilon$ values for each simulation. The simulated dataset is plotted in Fig. 5b,c for $\overline{\lambda}_{x,y}$ and Fig. 5d,e $\overline{\epsilon}_{x,y}$, respectively.
@@
The parameters extracted from these simulations – crosshatch wavelength $\overline{\lambda}$ and amplitude $\overline{\epsilon}$ – are calculated the same as our previous analysis (schematic shown in Fig. 1c). We used these parameters to train a Random Forest machine learning model, which predicted the dislocation density from the AFM-derived $\overline{\lambda}$ and $\overline{\epsilon}$ values. The best hyperparameters for the model were determined through cross-validation and yielded a test $R^{2}$ value of 0.976, indicating our model captures nearly all of the observed variance in predicting dislocation density. More details about the model can be found in the supplemental materials.
@@
I Supplemental Material
The supplemental material document contains further details regarding the machine learning model data curation, model selection, and training.
@@ Acknowledgements.
The authors would like to acknowledge support for this project from the Office of Naval Research award number N00014-22-1-2764.

### v2

@@ I Introduction
Figure 1: a Schematic of near-surface InAs quantum well heterostructure. b Exemplary 20x20 $\mu$m AFM image for a near-surface InAs quantum well sample JS629. c Schematic of crosshatch waveform model, adapted from [4]. d-f Plots of quantum well mobility as a function of the average crosshatch amplitude ($\overline{\epsilon}$) and wavelength ($\overline{\lambda}$).
Specifically, near-surface quantum well heterostructures within the last decade have been extensively studied for topological superconductivity [5, 6] and gate-tunable superconducting circuits for quantum information[7, 8, 9, 10, 11]. These studies utilize material stacks in which the quantum well region is placed close to the surface of the wafer. In this case, the active region of the wafer is well-separated from the initial mismatched interface by the growth of substantial metamorphic buffer layers [12, 13]. Compared to buried quantum well structures which are well known to have carrier scattering dominated by background and ionized impurities [14], modulation in the surface structure and surface dangling bonds have a much stronger impact on near-surface quantum well systems. As a result, understanding the surface strain state and strain fluctuations becomes of greater importance. Therefore, in these heterostructures it is of great significance to develop models with which we can quickly understand and predict how these materials will perform in cryogenic applications without the need for long fabrication, cooldown, and measurement cycles.
@@
Figure 2: Deconvolution of AFM image. a Averaged color map for vertical crosshatch channel of sample JS629. b and c are the vertical and horizontal channels of sample JS629 after threshold processing to define the peaks and valleys of the waveform. d Overlaid vertical and horizontal channels from our AFM image processes algorithm on top of the original rotated AFM image from Fig. 1b.
Figure 3: Extraction of $\lambda$. Histogram plots in a and b for the vertical and horizontal channels, respectively, for the exemplary AFM image of sample JS629, presented in Fig. 1a. The overlaid line in both plots is a Gaussian fit to the histogram from which we extract a center of mass value and a line width. These values are listed in the legend of this plot.
II AFM Crosshatch Extraction and Evaluation
@@
Once the AFM images are processed and we obtain the thresholded channel plots as shown in Figs. 2b and c, we then extract the widths of each line and gap between lines. These values are plotted in separate histogram plots for the vertical and horizontal channels, examples of which are shown in Figs. 3a and b. The histogram shows extracted values for the AFM image presented in Fig. 1b and corresponds to the de-convoluted crosshatch graphs presented in Fig. 2b and c. We then fit this histogram to a Gaussian distribution to extract a center of mass (CoM) value and the full-width at half max (FWHM). In the context of a real material evaluation, the CoM is related to the crosshatch density (residual strain state), whereas the FWHM value is then related to the fluctuations in this strain state.
@@
III Linear Regression Model
Figure 4: Linear regression model fitting. a Linear regression model fitting to the full dataset. The red line is the model fit. Calculated R2 and p-values are stated in the inset. b Linear regression model fitting with $\lambda_{x}$ and $\lambda_{y}$ removed from dataset. The red line is the model fit. c Calculated residuals for model 1. d Calculated residuals for model 2.
We then fit the variables from the compiled dataset across all 36 samples to a multiple linear regression model using the stats package within base R statistical software (ver 4.3.3) [15]. The results of two of these model fits are presented in Fig. 4. Figs. 4a,c are the regression model and calculated residuals for Model 1 and Figs. 4b,d are the same for Model 2. The difference between models is that for Model 2, the $\overline{\lambda}_{x}$ and $\overline{\lambda}_{y}$ variables are removed from the fit. The resulting linear regression equations for Model 1 and 2 are presented in the supplemental material. To generate the plots in Figs. 4a, b, the model $\overline{\lambda}_{MFP}$ value is calculated using the linear regression equations found in the supplmental material using the experimental dataset that is obtained through our computer vision algorithm. The $R^{2}$ and p-values for each model fit are presented in the insets in Figs. 4a and b.
@@ III Linear Regression Model
IV 2D Crosshatch Simulation
@@
Figure 5: Crosshatch surface modeling. a Simulated crosshatch surface for a starting density of $1.4\times 10^{9}/cm^{2}$. b and c present exemplary experimental (“original”) and the 2D simulated AFM images, respectively, for comparison. The simulated image presented here is the 2D representation of the 3D crosshatch surface presented in a. d and e: $\overline{\lambda}_{x}$ and $\overline{\lambda}_{y}$, respectively, as a function of dislocation density from our crosshatch simulations. f and g: $\overline{\epsilon}_{x}$ and $\overline{\epsilon}_{y}$ as a function of dislocation density from our crosshatch simulations. In d-g, the red lines represent the average value across the repeat simulations for each dislocation density while the grey shaded area represents the first standard deviation.
Our AFM simulation utilizes the model for surface crosshatch proposed by Andrews et al. in 2002 [3]. The original model solves Hooke’s law in one dimension for an arbitrary elastic medium. The model is initialized by defining some starting density of dislocations along the a 1D line, and then integrating the strain field caused by the dislocations to a specified distance away from this starting point. This results in us obtaining an expression for the displacement field of the elastic medium at an arbitrary distance away. We first expanded the model to include two dimensions to generate a surface crosshatch in InAs from a given dislocation density. This model randomly places dislocations within a 20x20 $\mu$m region based on an initial user-defined dislocation density. The model equations are presented below in Equations 1 and 2 for calculating the displacement field in x and y, respectively.
@@
V Machine Learning Analysis
When generating the simulated dataset, we simulate the surface 1.5 $\mu$m away from the misfit interface for each dislocation density, consistent with the thickness of the material stack used in our experiments. For simplicity, we assume all dislocations form at the interface with the InP substrate. The simulation was run for 1000 iterations per dislocation density, for 18 different densities ranging from $3\times 10^{7}/cm^{2}$ to $3\times 10^{9}/cm^{2}$. With this, we then extract the simulated $\lambda$ and $\epsilon$ values for each simulation. The simulated dataset is plotted in Fig. 5b,c for $\overline{\lambda}_{x,y}$ and Fig. 5d,e $\overline{\epsilon}_{x,y}$, respectively.
@@ V Machine Learning Analysis
The parameters extracted from the AFM simulations – crosshatch wavelength $\overline{\lambda}$ and amplitude $\overline{\epsilon}$ – are calculated the same as our previous analysis (schematic shown in Fig. 1c). We used these parameters to train a Random Forest machine learning model, which predicted the dislocation density from the AFM-derived $\overline{\lambda}$ and $\overline{\epsilon}$ values. The best hyperparameters for the model were determined through cross-validation and yielded a test $R^{2}$ value of 0.976, indicating our model captures nearly all of the observed variance in predicting dislocation density. More details about the model can be found in the supplemental materials.
@@ VI Conclusion
VII Experimental Details
Material Growth: InAs near-surface quantum wells are grown in a Varian Gen II molecular beam epitaxy (MBE) system on 50.8 mm InP (001) wafers. The InAs quantum well structures are grown on a strain-relaxed graded buffer layer with top-layer composition of In0.81Al0.19As. The quantum well is grown to be 4 nm thick and is separated from the wafer surface by a 10 nm thick In0.81Ga0.19As barrier layer. More details and specifics about this structure growth can be found in reference [12].
Atomic Force Microscopy Imaging: Atomic force microscopy imaging is done with a NanoMagnetics Instruments AFM operating in a dynamical tapping mode. Preliminary AFM image normalization and processing is done using the Gwyddion software package [16].
Quantum Hall Characterization: Quantum Hall transport is measured in an Oxford TelsatronPT 1.5 K He4 cryostat. The 2-in wafers are diced into roughly 5$\times$5 mm pieces which then are etched in Transene Type-D (Transene Company) aluminum etchant to remove the top Al layer. Samples are then cleaned in solvents, rinsed in de-ionized water, and then blow-dried with dry nitrogen. Measurements are done in a Van der Pauw wiring configuration.
VIII Supplemental Information
The supplemental material document contains further details regarding the linear regression model used and the machine learning model data curation, model selection, and training.
@@ Acknowledgements.
We thank Stefano Martiniani for fruitful discussions. The authors would like to acknowledge support for this project from the Office of Naval Research award number N00014-22-1-2764.